OpenStudy (lukecrayonz):

(2/y+2)-(y/2-y)=(y^2+4/y^2-4) (2/y+2)-(y/2-y)=(y^2+4/y^2-4) @Mathematics

jhonyy9 (jhonyy9):

2(y-2)+y(y+2) y2 +4 ------------ = -------- (y2 -4) (y2 -4) 2y-4+y2 +2y y2 +4 ------------- = ---------- (y+2)(y-2) (y-2)(y+2) 4y-4+y2 -y2 -4 --------------- = 0 (y-2)(y+2) 4y-8 ---------- = 0 (y-2)(y+2) 4(y-2) -------------- = 0 (y-2)(y+2) 4 ------- = 0 (y+2)